Please help this problem is making crazy! I picked this problem because I thought it wouldn't be so hard, but I'm having a hard time anyway. Could you give me a hand. I really really thank-you. If the perimeter of a rectangle is 93in than what is the L & W?

Answer by stanbon(75887)

If the perimeter of a rectangle is 93in than what is the L & W?
----------------------
Perimeter = 2(L+W)
2(L+W) = 93
L+W = 46.5 in.
L = 46.5 - W
W = W
==================
Cheers,
Stan H.